A source radiates sounds of 1000 Hz and 2000 Hz uniformly in all directions. At a distance of 100m from the source the intensity at both frequencies is the same. It follows that the pressure amplitude of the 1000-Hz sound at the distance of 100m from the source is the same as that of the 2000-Hz sound at a distance of?
1
Expert's answer
2020-10-05T10:57:15-0400

For sound waves the intensity is defined as


"I=\frac{1}{2}\rho A^2\omega ^2v" or "I=\frac{(\Delta p_0)^2}{2\rho v}"


if "I_{1000}=I_{2000}" "\to" "\Delta p_{1000}=\Delta p_{2000}".


Yes. The pressure amplitude of the 1000-Hz sound at the distance of 100m from the source is the same as that of the 2000-Hz sound at the same distance.